OpenStudy (anonymous):

Find the magnitude & direction of the vector JK given points J(-6,-4) & K(-10,-4).

OpenStudy (anonymous):

\[\text{magnitude} = \sqrt{(x_2 - x_1)^2 + (y_2-y_1)^2}\] \[\text{direction} = \tan^{-1} (\frac{y_2 - y_1}{x_2 - x_1})\]